OpenStudy (amy2012):

Kerry earned a total of $900 last year on his investments. If $7000 was invested at a certain rate of return and $9000 was invested in a fund with a rate that was 2% higher. Find the 2 rates of interest.

OpenStudy (wolf1728):

x = 7,000 * (1+r) y = 9,000 * (1+ r +2%) x + y -16,000 = 900